If a pro basketball player has a vertical leap of about 30 inches, what is his hang time? Use the hang-time function V=48T^2
PLEASE HELPPPP!!!!

Answer by Theo(13342)

V equals the vertical leap.
T equals the hang time.

The formula is V = 48T^2

V = 30 inches.

Formula becomes:

30 = 48T^2

divide both sides of this equation by 48 to get:

30/48 = T^2

take the square root of both sides of this equation to get:

T = sqrt(30/48)

Use your calculator to solve to get:

T = .790569415 seconds.